### v1

# Modified Axelrod Model Showing Opinion Convergence And Polarization In Realistic Scale-Free Networks

Abstract
Axelrod model is an opinion dynamics model such that each agent on a two-dimensional lattice has a finite number of possible opinions. The model is not realistic as all agents will eventually share the same opinion on every issue. Here we identify certain limitations inherent in the Axelrod model and address them by modifying some of the model’s rules. Specifically, by considering agents in a scale-free network and incorporating a more realistic interaction rule with continuous values representing agents’ opinions, our modified model offers a more comprehensive exploration of opinion dynamics. Computer simulation of our model shows a notable trend in opinion polarization on all features in the majority of reasonable simulation parameters. To mitigate this polarization, we introduce agents that work actively to reduce opinion differences. However, our findings indicate limited success in the approach. This research contributes to the understanding of opinion dynamics within society and highlights the nuanced complexities that arise when considering factors such as network structure and continuous opinion values. Our results prompt further exploration and open avenues for future investigations into effective methods of reducing opinion polarization.
Axelrod model, opinion dynamics, opinion polarization, scale-free network
I Introduction
In socio-physics, opinion dynamics of individuals or groups is modeled via interactions among them. By investigating the emergence and convergence of individual opinions into collective opinions, the socio-physical approach aims to shed light on the underlying mechanisms. In particular, sociophysicists wanted to understand the cause and evolution of opinion formation and dissemination through social influence and networks, the conditions that lead to consensus or polarization, and the factors governing the resulting distributions of opinions [1].
Numerous models of opinion dynamics have been proposed throughout the years. They include the Sznajd model [2, 3], Hegselmann-Krause model [4], Galam models [5], Axelrod model [6], Deffuant-Weisbuch model [7, 8] and their variants [9, 10, 11, 12, 13]. Our research here focuses on the Axelrod model [6]. In this model, individuals that we call agents from now on are located on a two-dimensional lattice. The opinion of each agent evolves by interacting with neighboring agents according to a fixed set of rules to be described in Sec. II. In particular, Axelrod found that after a sufficiently long time, all agents on the lattice tend to share the same traits for all features. That is, the traits of opinion (or opinion for short) in society converge [6]. Thus, this model fails to capture the fragmentation of cultures or political stances in numerous countries and societies in our contemporary world.
Various groups of social scientists have proposed reasons for this fragmentation in the real world [14, 15, 16, 17, 18]. For example, Ashraf and Galor [14] argued that genetic diversity, shaped by the migration of humans out of Africa, is a fundamental determinant of cultural heterogeneity. Dandekar et al. [15] tried to explain the polarization by opinion formation. Socio-physical models could offer a complementary approach to understanding the dynamics of opinion fragmentation. Different scholars have tried different simulation models attempting to capture the fragmentation behaviors in society. For instance, Chau et. al. proposed a model for opinion formation based on the social judgment theory extended from the Jager-Amblard model [19, 13]. That model incorporates assimilation and contrast, which influence agents’ opinion acceptance or rejection. More importantly, it shows opinion clustering around extremes and/or moderate. Furthermore, with slow variation of model parameters, Chau et al. observed first-order phase transition in the distribution of opinion [13].
Could the Axelrod model be modified to give fragmented opinions or to incorporate cultural diversity in the long time limit? Various authors have addressed this issue over the years with partial success. For example, MacCarron et al. [20] considered a more complicated interaction rule and analyzed the clustering behaviors. They showed that by decreasing a certain tunable parameter of the model known as the agreement threshold, agents on the lattice will no longer be homogeneous in the sense that they have more than one set of features in the long run. Along another line, Battiston et al. [21] proposed that the interactions in real society are carried out in multiple layers, which correspond to different interests or topics. By considering this multiplicity in their model, they qualitatively presented new dynamical behavior that leads to the emergence and stability of cultural diversity final state [21]. Worth noticing that both MacCarron et al. and Battiston et al. considered discrete feature traits. Along a different line, Amorim modified the Axelrod model by aligning it with Bourdieu’s theory on the relation between an individual’s taste and the social environment [22].
This paper introduces a socio-physical model of opinion dynamics that builds upon the Axelrod model, incorporating continuous attitude change among agents within a scale-free network. Through analysis, our model demonstrates that agents within society gradually develop strong opinions based on our evolutionary rules. These findings are consistent with observed social behavior and hence give novel insights into the factors influencing collective human opinions. In addition, we attempted to solve the polarization problem that is so urgently needed in our currently highly divided and politicized society by modifying our model. Specifically, we set a small portion of agents who actively try to reduce opinion differences. Nonetheless, this approach can only reduce opinion polarization slightly. This work is an extension of a short project done by XZ under the supervision of HFC, the teacher of the course “Data Analysis And Modeling In Physics” [23].
II The Axelrod Model
One of the most frequently used ways of measuring individual opinions in a society is to collect data from attitudinal surveys over a long period of time. It is instructive to model the attitude evolution of individuals and interest groups. One such model was proposed by Axelrod in 1997 [6]. In the Axelrod model, agents are placed on a $N\times N$ square lattice with closed boundary conditions. An agent is conveniently labeled by its position $(i,j)$ on the lattice. Each agent is characterized by the same fixed set of $F$ features =$\mathcal{F}=\{f_{1},f_{2},\cdots,f_{F}\}$. Each feature takes only values from a fixed set of traits. For simplicity, each $f_{i}\in\mathcal{F}$ is taken from the same set of $q$ traits labeled $\{0,1,\dots,q-1\}$. As a result, each agent has $F^{q}$ possible internal states. For example, one of the features could be the primary language spoken by the agent, and the corresponding set of traits is the set of all contemporary languages. Another possible feature is one’s religious belief. Each feature of an agent is randomly assigned in the beginning. Furthermore, the system evolves for several time steps as follows.
  1. 1.
    In each time step, an agent $(i,j)$ is randomly picked.
  2. 2.
    One of the nearest neighbors of $(i,j)$ is randomly chosen.
  3. 3.
    Denote the number of features that this pair of agents have in common by $F_{\text{comm}}$. And denote the features that the two agents are different by $\mathcal{F}_{\text{diff}}$.
  4. 4.
    With probability $P(\text{interaction})=F_{\text{comm}}/F$, we randomly pick a feature in $\mathcal{F}_{\text{diff}}$ and the agent $(i,j)$ then copies the trait of that feature from its chosen neighbor. In what follows, we call $F_{\text{comm}}/F$ the cultural similarity between two agents.
Figure 1: Typical time evolution of 100 agents on a $10\times 10$ lattice with $F=5$ and $q=10$ according to the Axelrod model [6]. Adjacent agents are connected by different line styles according to their cultural similarities $F_{\text{comm}}/F$ (solid for $F_{\text{comm}}/F\leq 20\%$, dashed-dotted for $40\%$, dotted for $60\%$, dashed for $80\%$ and no line for $100\%$. It shows the convergence of opinions over time.
The system evolves according to the above rules repeatedly and its eventual behavior is analyzed. Fig. 1 reproduces the opinion convergence behavior of the Axelrod model using parameters in their paper, namely, for $10\times 10$ nodes on a lattice with $F=5$ and $q=10$ [6].
Worth mentioning is that throughout the years, many opinion dynamics models have been proposed. In particular, Deffuant et al. put forward several agent-based models with continuous opinion change such that the involved agents have limited ranges of acceptance and non-commitment, where only the impact of assimilation is taken into account [8, 7].
III Deficiencies Of The Axelrod Model And Our Remedy
The Axelrod model is simple and useful in opinion dynamics. However, it has several deficiencies. We point them out and report ways to modify the model to make it more realistic below.
III.1 Network Topology
In the Axelrod model, agents form a two-dimensional lattice. This network topology is artificial. More importantly, it is not compatible with the fact that many social structures can be described by scale-free networks [24]. In simple terms, all but those agents in the lattice boundary in the Axelrod model have equal strength in influencing others or being influenced by others. In contrast, scale-free network topology allows the existence of key opinion leaders that are so commonly found in the real world. Because of these, we propose putting agents in an (undirected) scale-free network in which only neighboring agents are allowed to interact directly. For simplicity, we generate such a network using the Barabási–Albert (B-A) model [25] as it is one of the most studied and used scale-free networks. Besides, it has the additional advantage that it can be efficiently generated. In this study, we simplify matters by considering non-evolving networks. It is instructive to consider evolving networks in future studies.
III.2 Feature Traits
In the Axelrod model, every agent is described by a set of features, and each feature can have several possible discrete traits [6]. While there are features that are naturally described by discrete traits, the Axelrod model is incomplete as traits of a feature need not be discrete. For example, one of the features could be the agent’s fluency in a particular spoken language. Here, the possible traits are more naturally described by a continuous real number than discrete states. More importantly, using continuous traits is a natural platform to model the possibly increasingly polarized opinion distribution. We, therefore, modify the model by replacing the discrete set of traits with a real number representing the strength of the agent’s feature. We also call this type of continuous feature traits opinions of the agent.
How could we take care of the increasingly polarized opinion distribution? The naive way of representing opinions by a real number may not work too well owning to rounding errors and numerical instabilities in opinion evolution. Our resolution is to borrow the idea of the bounded confidence model [7, 4, 8, 19] by representing opinions using a real number in a bounded interval. Here we stress that our approach differs from the bounded confidence model in two ways. First, we restrict opinions in the open interval $(-1,1)$ instead of the closed one $[-1,1]$ so that there is always room for further opinion polarization. Second, we replace the standard Euclidean distance between two opinions $d_{\text{E}}(p,q)=|p-q|$ by the hyperbolic distance metric (of the Poincaré ball model), namely,
$d_{\text{h}}(p,q)=\ln\left(\frac{|1-pq|+|q-p|}{|1-pq|-|q-p|}\right)$ (1)
for any $p,q\in(-1,1)$. In other words, opinions are placed in a hyperbolic distance space. Note that $d_{\text{h}}$ diverges if and only if $(p,q)$ approaches $(1,-1)$ or $(-1,1)$. Hence, the hyperbolic distance metric on the interval $(-1,1)$ conveniently describes opinion polarization without significantly affecting the accuracy of our simulations. As for the initial values of each trait, instead of drawing from a uniform distribution in $(-1,1)$, we use a more realistic normal distribution with respect to the hyperbolic metric. Its detailed construction can be found in Sec. III.3 below.
III.3 Interaction Rules Among Agents
Recall that agents change their traits according to rule 4 in the Axelrod model. This rule can only increase the cultural similarity between the chosen pair of agents. More importantly, numerical experiments tell us that the system becomes homogeneous eventually under this rule [6]. In other words, rule 4 is too cohesive. Besides, it is not consistent with the observation that cultural diversity among groups of people appears even in apparently culturally homogeneous societies [14, 15, 16, 17].
We may also need to take into consideration the boomerang effect, which is the phenomenon in social interaction where attempts to persuade or restrict someone’s freedom of opinion result in the adoption of an opposing view. This effect occurs when persuasive communication happens with an unintended position, which often strengthens the original attitude an individual holds that it sought to alter. Hovland, Janis and Kelly first identified and named the boomerang effect in 1953 [26]. The boomerang effect has been observed in various contexts [27, 28], including consumer behavior, psychological interventions, health policies, and, relevant to our discussion here, political communication. In political contexts, attempts to correct misinformation through mock news articles sometimes lead to increased belief in inaccurate information among certain groups [29].
To study and model the positive persuasion as well as the boomerang effects in social interactions, we introduce a parameter $a\geq 0$ known as the agreement threshold. It is used to determine whether a social interaction results in opinion convergence or divergence. In each time step, we randomly pick an agent from the scale-free network plus one of its neighbors. We count the number of features in which the hyperbolic distance between the opinions of the two agents is smaller than the agreement threshold $a$. If this number is greater than or equal to, say, half of the total number of features, the two agents are said to be “similar”. Otherwise, they are “dissimilar”. We hypothesize that upon interaction, opinions of similar agents are likely to converge whereas those of dissimilar ones are likely to diverge. Besides, an agent’s persuasion power is proportional to the number of neighbors it has. This models the effect that a key opinion leader can change the opinions of his/her followers more effectively than an ordinary individual. Specifically, we state the details of our model below.
  1. 1.
    [Scale-free network generation] A B-A network with $N$ vertices is generated with each newly added vertex connecting to $m$ existing nodes using their standard preferential attachment mechanism [25]. Each vertex of the network represents an agent and only those agents joined by an edge in the network can communicate directly with each other.
  2. 2.
    [Opinion initialization] Each agent is characterized by the same finite set of $F$ features =$\mathcal{F}=\{f_{1},\cdots,f_{F}\}$. The opinion of agent $X$ on feature $f\in\mathcal{F}$, denoted by $X_{f}$ is randomly and independently drawn from $\mathcal{N}_{h}(\bar{x},\sigma)=\mathcal{N}_{h}(0,0.5)$, namely, the normal distribution with mean 00 and standard deviation $0.5$ in the hyperbolic distance metric. More precisely, if $r$ is the value of the random variable drawn from $\mathcal{N}_{h}(0,0.5)$, then the opinion is set to $v$ with $d_{\text{h}}(0,v)=r$. Besides, $v\geq 0$ if $r\geq 0$; and $v<0$ if $r<0$.
  3. 3.
    [Opinion evolution] An agent $X$ is randomly picked. Then, a neighbor of $X$, say $Y$, is randomly selected. A feature $f$ is randomly selected from $\mathcal{F}$.
    1. (a)
      [The case of similar agents] If the two agents are similar, namely, the hyperbolic distances between at least $\kappa=1/2$ of their opinions are less than the agreement threshold $a$, then $X_{f}$, the opinion of feature $f$ for agent $X$, changes to $\tilde{X}_{f}$ where
      $d_{\text{h}}(\tilde{X}_{f},X_{f})=\frac{\mu c_{Y}d_{\text{h}}(X_{f},Y_{f})}{c_{X}+c_{Y}}$ (2a)
      with $\operatorname{sgn}(\tilde{X}_{f}-X_{f})=\operatorname{sgn}(Y_{f}-X_{f})$. Here, $c_{X}$ and $c_{Y}$ are the degrees of nodes $X$ and $Y$, respectively. In addition, $\mu>0$ is a parameter modeling the effectiveness of persuasion. Likewise, $Y_{f}$ changes to $\tilde{Y}_{f}$ where
      $d_{\text{h}}(\tilde{Y}_{f},Y_{f})=\frac{\mu c_{X}d_{\text{h}}(X_{f},Y_{f})}{c_{X}+c_{Y}}$ (2b)
      and $\operatorname{sgn}(\tilde{Y}_{f}-Y_{f})=\operatorname{sgn}(X_{f}-Y_{f})$. In other words, the hyperbolic distance of the new opinions of the selected feature trait between the two agents decreases upon interaction.
    2. (b)
      [The case of dissimilar agents] If the two agents are dissimilar (that is, they are not similar), then we randomly pick a feature $f$ among those with $d_{\text{h}}(X_{f},Y_{f})\geq a$. The value of $X_{f}$ changes to the unique $\tilde{X}_{f}$ satisfying Eq. (2a) and $\operatorname{sgn}(\tilde{X}_{f}-X_{f})=\operatorname{sgn}(X_{f}-Y_{f})$. Similarly, $Y_{f}$ changes to the unique $\tilde{Y}_{f}$ satisfying Eq. (2b) and $\operatorname{sgn}(\tilde{Y}_{f}-Y_{f})=\operatorname{sgn}(Y_{f}-X_{f})$. Thus, the hyperbolic distance of the new opinions of the selected feature trait between the two agents increases upon interaction.
    The opinion evolution in Step 3 is repeated $Nt/2$ times, where $t$ is the average number of interactions an agent has with its neighbors.
Note that we use the probability distribution $\mathcal{N}_{h}(0,0.5)$ to initialize opinions to avoid opinion bias as well as to give an initial opinion diversity in our initial ensemble. Last but not least, we remark that our evolution rules do not change the nature of the agent pairs picked in Step 3 in the sense that they are similar (dissimilar) before and after the interaction. Nevertheless, the nature of each of the picked agents with its other neighbors may change after the interaction.
IV Simulation Results And Effects Of Different Model Parameters
IV.1 Methodology
We study the dynamics of our modified Axelrod model using Python. The scale-free network is created using the Networkx package and most of the analytical work is done using the NumPy package. Figures are produced by the Matplotlib package.
Parameter Description Default value
$N$ Number of agents 1000
$m$ Number of edges attached from new node to existing nodes while constructing the initial B-A network 4
$F$ Number of features 11
$t$ Mean number of interactions per agent 200
[…]
$a$ Agreement threshold that determines the similarity of agents 0.25
$\kappa$ Fraction of feature traits used to determine the similarity of agents 0.5
$\mu$ Persuasion effectiveness parameter 0.25
Table 1: Default values of different parameters used in our simulation.
Figure 2: Histograms showing the time evolution of opinion distribution with all but the parameter $t$ taken from Table 1. The $x$-axes of all subplots are the opinions in $(-1,1)$ and the $y$-axes are the occurrence frequencies of the opinion. The first six subplots have common $y$-axis limits, from 00 to $750$. Although the initial opinions concentrate around 00, the system evolves to a state of extreme opinion polarization in the long time limit in which nearly all agents have extreme opinions on all features.
We investigate the effect of each model parameter on the development of the system by varying one of them and fixing all the rest to their corresponding default values as tabulated in Table 1. We report our simulation results in histograms on the opinion distribution of feature traits. Since our initialization and evolution rules apply to different feature traits in the same way, opinion distributions of different feature traits averaged over a sufficiently large ensemble of independent runs should be the same. This is indeed what we have observed in our simulation. Consequently, to save running time, we lump the distributions of all $F$ feature traits of every agent in plotting the histogram. In each histogram, we divide the opinion into $50$ evenly spaced bins (in the usual Euclidean distance) each with a bin width of $0.04$. Each of the simulation results presented below is taken from a single run randomly selected over a sample of 10 runs. In fact, all runs give very similar distributions.
IV.2 Feature Evolution
We begin by investigating the time evolution of opinions by fixing all other model parameters to the ones listed in Table 1. Figure 2 shows the time evolution of agents opinions. Clearly, opinion spreads as time progresses. When the mean number of interactions per agent $t$ is less than $\approx 80$, the opinion distribution roughly follows the Gaussian distribution whose mean is zero and the standard deviation increases gradually. When $t\gtrsim 80$, extreme opinions start to emerge. In addition, the fractional increase of extreme opinions accelerates. Eventually, extreme opinions dominate, and nearly all agents have extreme opinions on all features. That is to say, opinions become extremely polarized. This result can be qualitatively understood as follows. Since the agreement threshold $a$ is less than the standard deviation of the initial opinion $\sigma$, a sizable fraction of neighboring agents is dissimilar. Their interactions gradually lead to opinion polarization. Once a non-negligible portion of the opinion is polarized, our evolution rules accelerate further polarization for it is harder and harder to find a similar agent pair which consists of an agent with moderate opinion and another agent with an extreme opinion on the same feature. As a consequence, it is no longer possible to pull enough extreme opinion agents back to moderate ones. Thus, an extremely polarized society is formed in the end. Readers will find in Secs. IV.3IV.5 that the formation of extremely polarized opinion is a genuine and common feature of our model using reasonable parameters rather than the result of a cherry-picked set of parameters. (Surely, opinions are never polarized if $\sigma\approx 0$ or $\kappa\approx 0$. However, these sets of parameters are atypical and unrealistic.)
Figure 3: Opinion distributions after $Nt/2$ time steps with $t=200$ for different values of $N$. These subplots show no significant difference in opinion distribution for different $N$.
Figure 4: Opinion distribution after $Nt/2$ time steps with $N=1000$ and $t=200$ for different values of $m$. These subplots show no significant difference in opinion distribution for different $m$.
Figure 5: Opinion distribution after $Nt/2$ time steps with $N=1000$ and $t=100$ for different values of $m$. Compared with Figure 5, these subplots show that longer equilibration times are needed to reach the extremely opinion depolarization state as $m$ increases.
IV.3 Effect Of The Number Of Agents
We now study the effect of the number of agents $N$ while fixing other parameters to their default values listed in Table 1. As shown in Figure 3, there is no significant difference in the distribution between different histograms even if $N$ varies over four orders of magnitude from $10^{2}$ to $10^{5}$. This result is not surprising because, upon coarse-graining, the resultant network is still scale-free. And since the formation of extreme opinion polarization is the generic behavior of the system, the effective parameters of the renormalized system likely fall within this generic region. Finally, we find that the convergence rate of the opinion polarization is almost independent of $N$. That is to say, for a given mean number of interactions per agent $t$, the distribution between different histograms is statistically the same regardless of $N$.
IV.4 Effect Of The Average Degrees Of Agents In The Scale-free Network
Here we investigate the effect of the average degree of the scale-free network $m$ on the system evolution by varying it from $4$ to $48$. As shown in Figure 4, there is no significant difference in the final steady state of the opinion distribution among all agents. They all evolve to a highly polarized opinion state.
To study the rate of convergence to this highly polarized state, we reduce the average number of interactions per agent $t$ from $200$ to $100$. Figure 5 depicts that the equilibration time to the extremely depolarized state increases as $m$ increases. This result can be understood as follows. An agent has more neighbors on average as $m$ increases. Consequently, the opinion of an agent is influenced by more neighbors each has statistically independent opinions on each feature trait initially. Thus, it takes a longer time for opinions of the system to polarize.
Figure 6: Opinion distribution after $Nt/2$ time steps for $N=1000$ and $t=200$ for different values of $a$. Subplots (a)–(h) depict that opinions of the equilibrated states are extremely polarized when $a\lesssim\sigma$. Subplots (j)–(l) show that the system reaches consensus of a single moderate opinion in the long run if $a\gg\sigma$. In between these two parameter ranges, both moderate and polarized opinions coexist in the long run as in subplot (i).
IV.5 Effect Of The Agreement Threshold
We now study the effect of agreement threshold $a$ on the system. When $a\lesssim\sigma$, as shown in subplots (a)–(h) in Figure 6, the system equilibrates to two dominating peaks at a polarized opinion as in the generic case. This is because the initial opinion distribution is sufficiently diverse to allow a sizable portion of neighboring dissimilar agents. Their interactions seed more extreme opinions. In contrast, when $a\gg\sigma$, the system evolves to a single peak of moderate opinion and the extreme opinions population is very small. (See subplots (j)–(l) in Figure 6.) The reason is that with a large value of $a$, almost all agent pairs are similar. This leads to opinion convergence for at least a majority portion of the population. For an intermediate value of $a$, as depicted in Figure 6(i), the coexistence of a non-negligible number of extreme and moderate opinions is observed.
In summary, we find that the value of $a/\sigma$ is the most important factor determining the asymptotic behavior of the system. It is natural to expect that $a<\sigma$ in human societies for otherwise the boomerang effect can hardly be observed. The implication of our modified Axelrod model is that societies will evolve toward extreme opinion polarization — a phenomenon in our contemporary society that leads to all kinds of troubles making most people unhappy and some to go violence.
Figure 7: Evolution of opinion distribution for different portion of sympathetic agents $p_{\text{s}}$ and mean number of interactions per agent $t$. All other parameters are set to default values listed in Table 1.
Figure 8: Evolution of opinion distribution for different values of $p_{\text{s}}$ after the system has reached a steady final state. All parameters taken are the same as those in Figure 7.
IV.6 Effects Of Other Parameters
Last but not least, we consider the effects of varying each of the following parameters — $F$, $\kappa$ and $\mu$. (There is no need to consider the variations of $\bar{x}$ and $\sigma$. This is because the opinion evolution rule in Step 3 of our model depends only on the hyperbolic distances between opinion pairs. Changing the value of $\bar{x}$ simply shifts the opinion distribution by $\bar{x}$ in the hyperbolic distance metric. Besides, rescaling $\sigma\mapsto s\sigma$ and $a\mapsto sa$ just multiply the hyperbolic distance between the resultant $X_{f}$’s by the same factor $s$. Thus Step 3 implies that the resultant opinion distributions are scaled by the factor $s$ in hyperbolic distance. These two changes do not affect the number of agents with moderate as well as extreme opinions.) We find that varying $\kappa$ alters the final opinion distribution while varying the others do not. This is consistent with the analysis above. The reason is that by varying $\kappa$, we increase the fraction of similar agents and hence slow down the emergence of polarized opinion. Therefore, the simulation result is very similar to the result shown in Figure 6. We do not show the corresponding histograms here to save space.
V Reducing Opinion Polarization By Introducing Sympathetic Agents
Could we make the opinion distribution of the eventual states less polarized within the framework of our modified Axelrod model? We would like to investigate possible interventions that could alleviate extreme opinion polarization. As a general principle, we would like to make our interventions as small as possible, hopefully in the form of making small changes in our rules. This is because changing human psychological behavior en masse is not easy, at least in the short term. Only these kinds of small changes have a real chance to occur in practice. Here we propose to alter the behavior of a small fraction of agents. Specifically, we fix a small portion $p_{\text{s}}$ of the agents to be sympathetic. A sympathetic agent always interacts with its neighbors using the rules of a similar agent. One way to think of it is that each agent has a different agreement threshold, and that of a sympathetic agent is arbitrarily large. Their readiness to be influenced and persuaded by all kinds of opinions is the key to reaching a consensus in the population.
Here, we investigate the opinion evolution of systems with sympathetic agents in two ways. Figure 7 depicts the opinion distribution evolution when sympathetic agents are assigned in the beginning. As expected, the higher the value of $p_{\text{s}}$, the smaller the deviation in the attitudes of all agents. Although a sizable number of extremists is still present, a significant fraction of moderates remain. Nevertheless, this effect is significant only when $p_{\text{s}}\gtrsim 15\%$. Such a high percentage of sympathetic agents is very hard if not impossible to achieve in reality.
A more pressing issue is how to reduce society polarization in an already polarized society. Therefore, we study what happens if we change a fixed portion of agents from “normal” to sympathetic after the society has evolved almost to its final state. More precisely, we first let the system evolve with $t=200$ using parameters listed in Table 1 so that it has almost equilibrated. Then, we will randomly pick and fix $p_{\text{s}}$ of the agents and make them sympathetic. Finally, we further evolve the system for an additional $t$ up to $400$. From Figure 8, we observe that one can significantly reduce opinion polarization only when $p_{\text{s}}\gtrsim 60\%$, a requirement that is even harder to achieve.
VI Conclusions
In this paper, we have identified deficiencies in the Axelrod model of opinion dynamics and proposed a modified one as an attempt to build a more sophisticated model to describe opinion dynamics in human society. Major modifications include using scale-free networks replacing the static two-dimensional lattice grid together with a new set of interaction rules for similar and dissimilar agents. Our simulation results show that over essentially all reasonable parameters agents’ opinions on all features eventually become more and more polarized. Finally, we tried to reduce this extreme opinion polarization by introducing sympathetic agents that are open to talking and actively interacting with everyone. Their presence reduces opinion polarization but only when they appear en masse.
Our model is flexible and can be modified to consider more realistic cases, even including evolution on the network. Our research is a small first step toward finding viable methods to understand opinion dynamics and reduce opinion polarization in our contemporary society.
Acknowledgements.
XZ was partially supported by The Innovation and Technology Scholarship of the Hong Kong Federation of Youth Groups.